Express the confidence interval ( 267.8 , 520.4 ) in the form of ¯ x ± M E .
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (267.8+520.4) / 2=394.1

Margin of error = E = Upper confidence interval -   = 520.4-394.1=126.3

in form of

¯ x ± M E

394.1 ±126.3